The energy required to increase the speed of a certain car from 19 m/s to 23 m/s is 185 kJ.

What is the mass of the car?

At what speed is the car’s kinetic energy equal to 185 kJ?
Answer #1

The mass will be

2E 2×185×103 f-U232-19 2202.38kg m=

The velocity will be

v = \sqrt{\frac{2E}{m}} = \sqrt{\frac{2\times185\times10^3}{2202.38}} = 12.96ms^{-1}